OpenStudy (anonymous):

Solve the following system by graphing and identify the point of intersection. -0.6x-0.5y=1.3, 0.5x-0.5y=-2

OpenStudy (anonymous):

first, do you know how to solve for y?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

what do you negative? I'm abit lost.

OpenStudy (anonymous):

I'm not sure either thats exactly how it was written

OpenStudy (anonymous):

to solve for y, take -0.6x to the other side you get -0.5y=1.3+0.6x now divide 1.3 and 0.6x by -0.5

OpenStudy (anonymous):

-2.6 and -1.2x ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have y=-2.6-1.2x. now graph it and see where it intercepts the x axis

jhonyy9 (jhonyy9):

like a first step for you dont need working with decimals in case of every two equations multiplie both sides by 10 so than will get - 6x -5y =13 5x-5y= -20 hope this will help you solving it more easy